Overall the City of Long Beach fared considerably well during the last two days of intense rainfall, and we’re preparing for more rain today. The Emergency Communications and Operations Center (ECOC) was again activated this morning at 8:00 am to manage the City’s response.
We experienced over 5 inches of rain in the last three days. Yesterday over 1.5 inches fell within a one-hour period, and on Tuesday, nearly one inch in a half-hour period. Our systems are designed to handle one inch of rain an hour. Any more than one inch in an hour overwhelms the flow of runoff from the streets into the LA River and San Gabriel River.
We expect today’s rainfall to be constant for several hours this afternoon, but not as intense as the powerful rainfall on Tuesday and Wednesday that caused much of our problems. Yesterday, the City of Long Beach appeared to bear the brunt of the storm, receiving more rainfall than most other areas in the region.
Here’s a brief rundown of the City’s activities:
- Handed out more than 500 tons of sand, and expect to have given out more than 80,000 sandbags in a 3-4 day period.
- Public Works crews responded to 240 flooded location on Wednesday.
- Fire crews responded to 36 water-related calls between 3:00 and 5:00 pm on Wednesday, including six submerged vehicles and one vessel assist.
- Beach Maintenance crews have removed 5,000 tons of debris.
- Beach Maintenance crews attending to berms at Juniper and the Peninsula.
- The City of Long Beach provided sand and bags to the Los Angeles County Sheriff’s Air Operations at Long Beach Airport.
Sand and Sandbags continue to be available to the public at five fire stations, and residents are encouraged to call 562.570.2726 to report weather-related issues such as flooded streets, clogged storm drains, downed trees, etc.
Here’s a brief status update:
The Main Library is closed today, due to flooding from the storm. Staff will be given the choice to take vacation, personal leave or be redeployed. Library staff will post signs, update the library's website and notify branch libraries.
The Nature Center is closed today, due to the poor condition of the trails, which are extremely muddy with a lot of standing water. School tours have been cancelled for the remainder of the week.
3 rd District – 8 th and Bennet and the Roswell area are fine. Three Dumpsters have been provided to each of these areas.
University Park Estates (back of 6 th St.) Opened flap gate. Waters flowed better.
4 th District - Termino (10th to Anaheim) Remains closed.
5 th District - Spring Street Tunnel (Lakewood to Kilroy) has been reopened.
Lakewood @ 405 has been reopened.
- Investigating report of a downed electrical wire at E Monlaco Rd & N Charlemagne Ave. SCE notified.
6 th District – 25 th and Pine is closed, but we are reviewing the need to keep closed.
7th District - WB Willow @ TI Freeway, Intersection closed but stormdrain clear.
- Arlington Ave. area is doing fine.
Santa Fe and the 405 is open.